# Accurately simulating core-collapse self-interacting dark matter halos

### v2

@@ Appendix C Effective cross-section of the isolated halo
In this appendix we discuss the differences between the velocity-independent and dependent runs for the isolated halo as shown in Sect. 3.8. In addition to Fig. 12, we display the mass enclosed within 0.15 kpc as a function of time in Fig. 18. We choose a larger radius to reduce the uncertainty in the measured mass. Importantly, we rescale the time axis of the velocity-independent runs to match the collapse phase of the velocity-dependent simulations. It becomes visible that there is a qualitative difference between the simulations, given that the core-expansion phase cannot be matched simultaneously. We note, that the collapse time scale of the halo does not scale inversely proportional with the cross-section. This is because the halo is partially in the short-mean-free-path regime. If we nevertheless assume this scaling, the rescaled version of the velocity-independent cases would correspond to a cross-section of $\sigma_{\mathrm{V}}/m=49.6\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$.
Figure 18: Enclosed mass and energy conservation as a function of time. We show the same simulations as in Fig. 12 (O, P, R, and S in Tab. 1). The upper panel gives the mass within a radius of 0.15 kpc and the lower panels shows how well the total energy is conserved. The time for the simulations with a velocity-independent cross-section is rescaled to match the collapse phase of the simulations with velocity-dependent interactions.
The effective cross-section (Yang & Yu 2022), $\sigma_{\mathrm{eff}}$ (see Eq. (11)) is a common way to map the effect of SIDM models with different velocity and angular dependencies on each other (see also Yang et al. 2023b). It can be expressed as
@@
To allow for a better understanding of the difference between velocity-dependent and velocity-independent cases, we compute the effective cross-section (Eq. (11)) as a function of radius. Here, we do not use a single characteristic velocity dispersion, but instead use the value of the velocity dispersion at each radius from the simulation. In Fig. 19, we show the results for the simulation with isotropic scattering (run R according to Tab. 1). We can see that the effective cross-section initially is large at small and large radii but low for the intermediate range. At the early stage of core expansion, the constant cross-section $80\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ is a reasonable approximation for the effective cross-section. However, at later stages, the effective cross-section decreases a lot, which is a consequence of the increasing velocity dispersion and is in line with the long collapse time that we find in Fig. 12. This continuous decrease during the collapse may limit the usability of the effective cross-section using a single characteristic velocity dispersion to map the effect that different SIDM models have on the evolution of an isolated halo on each other. For the velocity-dependent models, we consider $w=20\,\mathrm{km}\,\mathrm{s}^{-1}$, which is comparable to the characteristic velocity dispersion of the initial halo $17.28\,\mathrm{km}\,\mathrm{s}^{-1}$. Thus, the effective cross-section is sensitive to the change of the velocity dispersion in the halo during the core collapse. For models where $w$ is larger than the characteristic velocity dispersion, we expect that the mapping approach should work better. We will leave detailed studies on this topic for future work.
Figure 19: Effective cross-section as a function of radius. For the isolated halo evolved with an isotropic velocity-independent cross-section (simulation R according to Tab. 1), we show the effective cross-section for several times as indicated in the legend. In addition, we display the cross-section for the velocity-independent cases (black line).
@@ Appendix D Fit with King model
Figure 20: Density as a function of radius. Simulation results and a fit using the King model (Eq. (9)) are shown for different times of the late collapse phase. The results are for the isolated simulation with a resolution of $N=5\times 10^{7}$ particles (simulation Q of Tab. 1).
Figure 21: Parameters of fitted King model and further quantities as a function of time. The parameters that we determine by fitting a King model (Eq. (9)) to our highest-resolved simulation (Q in Tab 1) are shown in the top panel as a function of time. The bottom panel gives the mass enclosed in $r_{\mathrm{c}}$ according to the fitted model as well as the maximum circular velocity directly computed from the simulation data.
In this Appendix, we fit our highest-resolved simulation for the halo in isolation (simulation Q according to Tab. 1) with the King model (Eq. (9)). As for the satellite halo in Sect. 4.4, we use a limited radial range only, $r<0.5\,\mathrm{kpc}$.
In Fig. 20, we give the density profile together with the fitted King model. The King model appears to provide a reasonable description of the central region of a collapsing SIDM halo. However, it does not describe the outer regions well and also may eventually fail for the inner region at even later stages. Nevertheless it may work well enough to fit observational data as discussed by Zhang et al. (2025). We note that the central density from the fit, $\rho_{0}$, rises super-exponentially with time and at the same time, the mass within $r_{\mathrm{c}}$ is decreasing as a function of time as we show in more detail below.
@@ Appendix E Orbit of the satellite and tidal evolution
Figure 22: Trajectory of the satellite. The three-dimensional position of the subhalo for a CDM and an SIDM simulation is shown. The corresponding simulations are T and W according to Tab. 1, where their parameters are given.
Figure 23: Tidal radius and its enclosed mass as a function of time. For the satellite halo, evolved with different DM models, the tidal radius (solid lines) according to Eq. (12) and the mass within it (dashed lines) are shown. All simulations shown here share a resolution of $N=10^{7}$ particles and a gravitational softening length of $\epsilon=1.0\,\mathrm{kpc}$. Further information on the simulations T, W, X, Y, and Z is provided in Tab. 1.
@@
Next, we investigate the tidal evolution of the satellite for the DM models we have simulated. Here, we compute the tidal radius, $r_{\mathrm{tidal}}$, of the satellite halo. For this we implicitly define it as
@@
In Fig. 23, we display the tidal radius as a function of time for various DM models and show its enclosed mass. It is visible that all DM models evolve fairly similarly, especially the SIDM models. This is related to the fact that the tidal radius is typically relatively larger compared to the initial scale radius of the satellite halo ($r_{\mathrm{s}}=1.28\,\mathrm{kpc}$).
@@ Appendix F Effective cross-section of the satellite
Figure 24: The effective cross-section as a function of time. We compute the effective cross-section for different radial ranges from the average velocity dispersion using the satellite simulation for an isotropic velocity-dependent cross-section (simulation Y of Tab. 1). Moreover, we show the velocity-independent case, which does not have any radial dependence.
In this appendix, we study the effective cross-section given by Eq. (11) for our simulation of a satellite halo begin subject to an isotropic velocity-dependent cross-section. We compute the results given in Fig. 24 using simulation Y as specified in Tab. 1. The effective cross-section requires a characteristic velocity dispersion. Here, we use the average velocity dispersion within two different radial ranges, namely $r<0.05\,\mathrm{kpc}$ and $0.7\,\mathrm{kpc}<r<1.5\,\mathrm{kpc}$. In addition, we show the case of the constant cross-section of $80\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ as we use it for the simulations for a velocity-independent cross-section.
It is visible in Fig. 24, that the effective cross-section for the velocity-dependent case is larger in the centre of the halo (red, solid) than for the constant case (blue). This leads to an enhanced core formation. However, later on, the effective cross-section drops below $80\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. At about the first pericentre passage ($\approx 1.5\,\mathrm{Gyr}$) tidal stripping is effective and reduces the velocity dispersion of the satellite halo as discussed in Sect. 4.3. As a consequence, the effective cross-section sharply rises for the velocity-dependent case. This is the case at small (dotted) and large (dashed) radii. Subsequently, the effective cross-section at small radii starts to decline because of the increase in velocity dispersion due to the gravothermal collapse. At larger radii, the impact of the second pericentre passage is visible, increasing the effective cross-section for later stages.
@@ Appendix G Highest-resolved simulation
Figure 25: Density, velocity dispersion and Knudsen number as a function of radius. We show the results for our highest-resolved halo (simulation Q of Tab. 1) at various stages of the gravothermal evolution.
In this appendix, we provide additional information for our highest-resolved simulation evolved in isolation (simulation Q in Tab. 1). Figure 25 gives the density (top panel), the one-dimensional velocity dispersion (middle panel), and the Knudsen number (bottom panel) as a function of radius for several times. Here, the gravothermal evolution driven by a velocity-independent cross-section including the core expansion and collapse phase is well visible.

### v4

@@ Appendix C Effective cross-section of the isolated halo
In this appendix, we discuss the differences between the velocity-independent and dependent runs for the isolated halo, as shown in Sect. 3.8. In addition to Fig. 12, we display the mass enclosed within 0.15 kpc as a function of time in Fig. 18. We choose a larger radius to reduce the uncertainty in the measured mass. Importantly, we rescale the time axis of the velocity-independent runs to match the collapse phase of the velocity-dependent simulations. It becomes visible that there is a qualitative difference between the simulations, given that the core-expansion phase cannot be matched simultaneously. We note that the collapse timescale of the halo does not scale inversely proportional with the cross-section. This is because the halo is partially in the short-mean-free-path regime. If we nevertheless assume this scaling, the rescaled version of the velocity-independent cases would correspond to a cross-section of $\sigma_{\mathrm{V}}/m=49.6\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$.
Figure 18: Enclosed mass and energy conservation as a function of time. We show the same simulations as in Fig. 12 (O, P, R, and S in Table 1). The upper panel gives the mass within a radius of 0.15 kpc and the lower panels shows how well the total energy is conserved. The time for the simulations with a velocity-independent cross-section is rescaled to match the collapse phase of the simulations with velocity-dependent interactions.
The effective cross-section (Yang & Yu 2022), $\sigma_{\mathrm{eff}}$ (see Eq. (11)) is a common way to map the effect of SIDM models with different velocity and angular dependences on each other (see also Yang et al. 2023b). It can be expressed as
@@
To allow for a better understanding of the difference between velocity-dependent and velocity-independent cases, we compute the effective cross-section (Eq. (11)) as a function of radius. Here, we do not use a single characteristic velocity dispersion, but instead use the value of the velocity dispersion at each radius from the simulation. In Fig. 19, we show the results for the simulation with isotropic scattering (run R according to Table 1). We can see that the effective cross-section initially is large at small and large radii but low for the intermediate range. At the early stage of core expansion, the constant cross-section $80\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ is a reasonable approximation for the effective cross-section. However, in later stages, the effective cross-section decreases a lot, which is a consequence of the increasing velocity dispersion and is in line with the long collapse time that we find in Fig. 12. This continuous decrease during the collapse may limit the usability of the effective cross-section using a single characteristic velocity dispersion to map the effect that different SIDM models have on the evolution of an isolated halo on each other. For the velocity-dependent models, we consider $w=20\,\mathrm{km}\,\mathrm{s}^{-1}$, which is comparable to the characteristic velocity dispersion of the initial halo $17.28\,\mathrm{km}\,\mathrm{s}^{-1}$. Thus, the effective cross-section is sensitive to the change of the velocity dispersion in the halo during the core collapse. For models where $w$ is larger than the characteristic velocity dispersion, we expect that the mapping approach should work better. We will leave detailed studies on this topic for future work.
Figure 19: Effective cross-section as a function of radius. For the isolated halo evolved with an isotropic velocity-independent cross-section (simulation R according to Table 1), we show the effective cross-section for several times as indicated in the legend. In addition, we display the cross-section for the velocity-independent cases (black line).
@@ Appendix D Fit with King model
Figure 20: Density as a function of radius. Simulation results and a fit using the King model (Eq. (9)) are shown for different times of the late collapse phase. The results are for the isolated simulation with a resolution of $N=5\times 10^{7}$ particles (simulation Q of Table 1).
Figure 21: Parameters of fitted King model and further quantities as a function of time. The parameters that we determine by fitting a King model (Eq. (9)) to our highest-resolved simulation (Q in Table 1) are shown in the top panel as a function of time. The bottom panel gives the mass enclosed in $r_{\mathrm{c}}$ according to the fitted model as well as the maximum circular velocity directly computed from the simulation data.
In this appendix, we fit our highest-resolved simulation for the halo in isolation (simulation Q according to Table 1) with the King model (Eq. (9)). As for the satellite halo in Sect. 4.4, we use a limited radial range only, $r<0.5\,\mathrm{kpc}$.
In Fig. 20, we give the density profile together with the fitted King model. The King model appears to provide a reasonable description of the central region of a collapsing SIDM halo. However, it does not describe the outer regions well and may also eventually fail for the inner region at even later stages. Nevertheless, it may work well enough to fit observational data as discussed by Zhang et al. (2025). We note that the central density of the fit, $\rho_{0}$, rises super-exponentially with time, and at the same time, the mass within $r_{\mathrm{c}}$ is decreasing as a function of time as we show in more detail below.
@@ Appendix E Orbit of the satellite and tidal evolution
Figure 22: Trajectory of the satellite. The three-dimensional position of the subhalo for a CDM and an SIDM simulation is shown. The corresponding simulations are T and W according to Table 1, where their parameters are given.
Figure 23: Tidal radius and its enclosed mass as a function of time. For the satellite halo, evolved with different DM models, the tidal radius (solid lines) according to Eq. (12) and the mass within it (dashed lines) are shown. All simulations shown here share a resolution of $N=10^{7}$ particles and a gravitational softening length of $\epsilon=1.0\,\mathrm{kpc}$. Further information on the simulations T, W, X, Y, and Z is provided in Table 1.
@@
Next, we investigate the tidal evolution of the satellite for the DM models that we have simulated. Here, we compute the tidal radius, $r_{\mathrm{tidal}}$, of the satellite halo. For this, we implicitly define it as
@@
In Fig. 23, we display the tidal radius as a function of time for various DM models and show its enclosed mass. It is visible that all DM models evolve fairly similarly, especially the SIDM models. This is related to the fact that the tidal radius is typically relatively large compared to the initial scale radius of the satellite halo ($r_{\mathrm{s}}=1.28\,\mathrm{kpc}$).
@@ Appendix F Effective cross-section of the satellite
Figure 24: Effective cross-section as a function of time. We compute the effective cross-section for different radial ranges from the average velocity dispersion using the satellite simulation for an isotropic velocity-dependent cross-section (simulation Y of Table 1). Moreover, we show the velocity-independent case, which does not have any radial dependence.
In this appendix, we study the effective cross-section given by Eq. (11) for our simulation of a satellite halo begin subject to an isotropic velocity-dependent cross-section. We compute the results given in Fig. 24 using simulation Y as specified in Table 1. The effective cross-section requires a characteristic velocity dispersion. Here, we use the average velocity dispersion within two different radial ranges, namely $r<0.05\,\mathrm{kpc}$ and $0.7\,\mathrm{kpc}<r<1.5\,\mathrm{kpc}$. In addition, we show the case of the constant cross-section of $80\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$ as we use it for the simulations for a velocity-independent cross-section.
It is visible in Fig. 24, that the effective cross-section for the velocity-dependent case is larger in the centre of the halo (red, solid) than for the constant case (blue). This leads to an enhanced core formation. However, later on, the effective cross-section drops below $80\,\mathrm{cm}^{2}\,\mathrm{g}^{-1}$. At about the first pericentre passage ($\approx 1.5\,\mathrm{Gyr}$) tidal stripping is effective and reduces the velocity dispersion of the satellite halo, as discussed in Sect. 4.3. As a consequence, the effective cross-section sharply rises for the velocity-dependent case. This is the case at small (dotted) and large (dashed) radii. Subsequently, the effective cross-section at small radii starts to decline because of the increase in velocity dispersion due to the gravothermal collapse. At larger radii, the impact of the second pericentre passage is visible, increasing the effective cross-section for later stages.
@@ Appendix G Highest-resolved simulation
Figure 25: Density, velocity dispersion, and Knudsen number as a function of radius. We show the results for our most highly resolved halo (simulation Q of Table 1) at various stages of the gravothermal evolution.
In this Appendix we provide additional information for our highest-resolved simulation evolved in isolation (simulation Q in Table 1). Figure 25 gives the density (top panel), the one-dimensional velocity dispersion (middle panel), and the Knudsen number (bottom panel) as a function of radius for several times. Here, the gravothermal evolution driven by a velocity-independent cross-section including the core expansion and collapse phase is well visible.